OpenStudy (anonymous):

Solve x + 2y + z = 5 2x – y + z = 4 3x + y + 4z = 1

OpenStudy (anonymous):

add 2nd and third equation

OpenStudy (anonymous):

Add the 2nd and the 3rd equation first and tell what you got?

OpenStudy (anonymous):

you will get 5x+5z=5 x+z=1

OpenStudy (anonymous):

2x - y + z = 4 3x + y + 4z = 1 Just add them..

OpenStudy (anonymous):